# Humanoid Whole-Body Badminton via Multi-Stage Reinforcement Learning

### v1

@@ IV-B1 Generate Shuttlecock Trajectory for Training
To generate badminton flight trajectories for training purposes, we employed a physics-based simulation approach. This method adheres to the badminton dynamics model in [23], with the shuttlecock’s flight state updated according to the following equation:
@@
where $m$ represents the shuttlecock mass, $v$ denotes its velocity, $g$ is gravitational acceleration, and $L$ is the aerodynamic length, defined as $L=2m/\rho SC_{D}$. Here, $\rho$ is air density, $S$ is the cross-sectional area of the shuttlecock, and $C_{D}$ is the drag coefficient. The computed aerodynamic length $L$ used in our work is 3.4.
The generated shuttlecock flight trajectories were required to meet specific conditions to ensure data validity and relevance. To simulate realistic badminton hitting scenarios, we defined a hitting zone constrained by $x\in[-0.8,0.8]$, $y\in[-1,0.2]$, and $z\in[1.5,1.6]$, with a minimum traversal time of 0.8 seconds. This filtering process effectively eliminated trajectories that did not conform to realistic hitting requirements, thereby ensuring the quality of the training data. Hitting zone on $y$ direction is unsymmetrical due to right hand racket setting.
Finally, the selected trajectories were combined with corresponding traversal point data to construct a comprehensive training dataset containing position, orientation (tangential line of the trajectory) and timing features. These data are stored as tensors and can be directly utilized for subsequent training.
@@ IV-B2 Shuttlecock Trajectory Prediction for Deployment
In the deployment phase, an EKF-based trajectory prediction algorithm was developed to achieve real-time shuttlecock flight estimation and hitting point prediction. This approach accounts for aerodynamic drag, gravitational acceleration, and sensor noise, providing accurate target for the robot’s striking motion planning.
Our Extended Kalman Filter implementation adheres to the same badminton dynamics model as that used for shuttlecock trajectory generation. The system state vector is defined as:
$\mathbf{x}=[x,y,z,v_{x},v_{y},v_{z}]^{\mathsf{T}}$ (6)
where position [x,y,z][x,y,z]and velocity $v_{x},v_{y},v_{z}]$are expressed in the global coordinate frame.
The Extended Kalman Filter operates through two distinct phases: measurement update and prediction. When the predicted height first falls below the predefined hitting plane at $z=z_{target}$, linear interpolation is employed to compute the spatial coordinates and time of traversal. This calculated point is designated as the predicted hitting point, with its position and time expressed as: $\mathbf{p}_{\mathrm{ee}}^{*}=[x_{\mathrm{hit}},\,y_{\mathrm{hit}},\,z_{\mathrm{target}}]^{\mathsf{T}}$ and $t^{*}$.
The corresponding velocity vector at $t_{hit}$ is then converted into a quaternion representation to guide the robot’s end-effector orientation during the hitting motion.
@@ IV-C Prediction-free variant
We additionally train a prediction-free policy whose only change from the main pipeline lies in the observation space; the reward function, action space, training schedule (three-stage) remain identical. In this variant, the actor no longer receives the commanded hitting position, orientation and time $(p_{ee}^{*},q_{ee}^{*},t^{*})$. Instead, it ingests a sliding window of world-frame shuttle positions comprising the current frame and the previous five frames sampled at 50 Hz (i.e., $\{p_{\mathrm{sh}}(t),p_{\mathrm{sh}}(t{-}0.02),\ldots,p_{\mathrm{sh}}(t{-}0.10)\}$). From this short history the actor must implicitly infer the intended interception pose and timing. The critic retains privileged access to the ”actual” interception point. This design is enabled by how training data are generated: for each commanded target, we do not only sample the single interception point but also store the entire shuttle trajectory integrated by forward badminton dynamics. During training, the reward and the critic always know the ground-truth hit specification associated with the current shot, while the actor only observes the raw shuttle-history window.
A practical advantage is that deployment no longer requires a hand-tuned predictor (e.g., EKF plus aerodynamic model) to deliver $(p_{ee}^{*},q_{ee}^{*},t^{*})$. Instead, the controller conditions on directly measured shuttle positions, greatly reducing sensitivity to mismatches in physical or estimator parameters. To further improve robustness, we randomize aerodynamic parameters per shot, effectively treating the simulator as having an “aerodynamic patch” whose exact coefficients need not be identified. Intuitively, this mirrors human play: players do not know a specific shuttle’s drag constants a priori but can infer landing tendencies from a brief flight segment. In simulation we observe that the prediction-free policy achieves hitting with a small performance drop relative to the known-target setting, while simplifying deployment by removing the need to retune predictor and physics parameters. Real-robot validation of this variant is left to future work.
@@ V EXPERIMENTS
We evaluate our approach with five experiments that together probe accuracy, agility, robustness, and deployability: (1) a simulation Two-Robot Rally that measures sustained rally length; (2) a simulation Target-Known vs. Prediction-Free Comparison that contrasts two observation designs under matched conditions; (3) an EKF Prediction Accuracy study that quantifies trajectory and intercept-point errors; (4) a real-world Virtual-Target Swinging test that assesses swing speed and pose accuracy without a real shuttle; and (5) a real-world Ball-Machine Hitting test that measures outgoing shuttle speed, reachable interception range, and return landing distribution.
@@ V-B Simulation Results
Figure 3: Simulation. Figure (a) illustrates the Two-Robot Rally scenario, where two identical humanoid robots sustain a rally of 21 consecutive returns. Figure (b) demonstrates the Prediction-Free policy: the robot infers the optimal impact position and orientation solely from the first five recorded shuttlecock positions after serving. Figure (c) presents the Target-Known policy, where a predetermined hitting position is provided. The red sphere indicates the designated hitting location, while the green sphere confirms successful impact execution by the robot.
@@ V-B1 Two-Robot Rally
Each policy executes twenty hits sampled randomly. We report two quantities: the position and orientation errors at impact and the executed swing speed. As shown in Fig. 4, explicit target information is not indispensable, the policy can infer both the hitting target and the timing on its own with very small performance drop.
Figure 4: Policy comparison. The top part of this figure shows the position error for both strategies. The middle section of the figure shows the orientation error comparison, the orientation corresponds to the normal direction of the racket face. The bottom part of the figure compares swing velocity.
@@ V-D Real Robot Deployment
In our real world evaluation, we decompose the shuttlecock hitting task into two prerequisites, trajectory prediction and swing execution, followed by a final integrative test.
@@ V-D1 EKF Prediction Accuracy study
Prediction error refers to the deviation between the model’s predicted shuttlecock trajectory and the actual measured trajectory. Accurate trajectory prediction is critical for the robot to hit the shuttlecock at the right time and position. To quantify prediction errors, we collected 20 authentic badminton flight trajectories using a motion capture system. Partial segments of these trajectories were used as measurement input for the EKF. The EKF was then employed to generate predictions, which were subsequently compared against the ground-truth contact positions and timing. The comparative results are presented in Fig. 5. At 0.6 s before impact, the mean predicted position error falls below 100 mm, with the error bound sharply converging over time. By 0.3 s prior to contact, the mean position error has narrowed to 10 mm. Meanwhile, interception timing prediction achieves a mean error of approximately 20 ms at 0.6 s before hitting-already meeting task requirement—and similarly exhibits rapid convergence.
@@ V-D2 Virtual-Target Swinging
The racket swing error refers to the spatial deviation between the racket’s contact point with the shuttlecock and the sweet spot during execution. To quantify the robot swing error, massive randomly sampled target position was designated, and Mocap was employed to measure the center position of the racket face. The robot was subsequently instructed to execute the hitting motion, and the actual measured interception point was compared against the commanded target position, the corresponding results are presented in Fig. 6.
@@
We conducted a total of 71 virtual target swinging tests. The results showed a mean Euclidean distance error of 24.29 mm with an average racket speed of 5.32 m/s at impact. These results verify that the robot can effectively control the contact point within the radius of the racket face during complex hitting tasks, confirming its practical feasibility for performing high-precision swinging actions.
@@ V-D3 Real-World Shuttle Hitting
When integrating two systems together in the real world, we set a ball machine to serve the shuttlecock, adjust its’ strength and pitch-yaw angle, to examine the capability of the policy. Due to the shuttle’s flat slope being limited in indoor Mocap environment, which left a very short window to react, we constraint the interception area in a reasonable range. We collect 25 hitting results; except for one hitting on the edge of racket frame, the robot all achieves successful shuttle return with coordinated and agile motion. As shown in Fig. 7, the robot intercepts the shuttlecock above the ground at about 1.5 m with the range of 98 cm x 42 cm. This is a relatively large area for a robot height only 1.28 m. We also observe that the robot not only manage to intercept the ball but also return it with high quality, resulting in a sharp slope demonstrated in Fig. LABEL:fig:figure1, with the highest trajectory point over 3 m. Moreover, we measure the return shot landing distance, except for 3 times hitting the ceiling and 1 time hitting on the racket edge it achieves 3.5 m on average away from the interception area.
Figure 7: Real-World Shuttle Hitting. This figure captures the robot’s actual hitting postures at the two opposing boundaries of the 98 cm × 42 cm interception area.
@@ VI-A System Design
We train a single whole-body policy that coordinates stepping and striking rather than layering an arm controller on top of a lower-body controller. In badminton, hit quality depends on how the legs reposition and supports the arm’s exertion while maintaining balance. The subtle movements of the lower limbs can effectively adjust the posture of the racket.
Our multi-stage curriculum proceeds in three clear steps: learn to reach interception regions with stable footwork; add a sparse, high-value hit signal so a full swing emerges and follow a sigma curriculum to improve accuracy with light style terms; finally remove most lower-limb regularizers to avoid gradients fighting the true objective of precise, energy-efficient hitting. After this removal, hitting metrics continue to improve while energy and torque costs drop, indicating progress toward task-optimal objective. Coupling position and orientation in the hit reward further prevents the policy from over-optimizing one at the expense of the other.
@@ VI-B Limitations and Future Work
Dependence on the Mocap arena with limited ceiling height constrains shuttle trajectories to lower arcs driven by a ball machine. During testing, the shuttle occasionally hit the ceiling, and the limited flight range makes long multi-ball rallies with human partners difficult. Higher ceilings and broader interception bands should reveal more of the controller’s potential.
We plan to deploy the prediction-free variant on hardware. It removes explicit target prediction at test time but requires a strictly timed 50 Hz history of shuttle positions. We will add robust buffering and time-stamping so the actor can infer velocities and timing from consistent spacings.
Moving toward pure vision will require reliable visual odometry and policies that keep the shuttle within FOV during aggressive motion. A head-mounted camera that actively tracks the shuttle, together with learning signals that encourage aligning the view with the shuttle trajectory, offers a practical path to remove external tracking.
Above the low-level controller, we will train a higher-level policy by self-play to decide where to intercept, when to swing, and what racket-face orientation to command, then invoke the whole-body controller as a motor primitive. Current failures arise when consecutive targets are too far apart for the available maneuver time. We will address this with wider target sampling and reward shaping that explicitly encourages high-speed legged maneuvers.
Finally, this framework provides a promising path to adapt the pipeline to tennis or squash by replacing the shuttlecock dynamic “patch” and adjusting interception style, with modest reward retuning to reflect different sport-specific behaviors.
@@ VII Conclusions
In this work, we present the first real-world humanoid badminton system powered by a unified whole-body RL policy trained with a three-stage curriculum. The curriculum progresses from guiding lower-body footwork, to inducing a full swing and gradually tightening positional and orientational accuracy, and finally removing most gait regularizers to focus learning on precise, robust hitting. The policy produces human-like badminton behaviors that coordinate stepping and swinging under tight time windows, without any expert demonstrations.
In simulation, two identical humanoids sustain rallies with a best run of 21 consecutive returns, achieving a hitting accuracy within 0.10 m position error and 0.2 rad orientation error. A separate comparison shows that the prediction-free variant matches the counterpart behavior that uses known hitting targets, while remaining robust to moderate aerodynamic variation. On hardware, the controller achieves swing speeds of about 5.3 m/s and returns ball-machine serves with outgoing shuttle speeds up to 10 m/s under sub-second timing, transferring zero-shot through domain randomization. We believe this framework marks an important step toward agile, interactive whole-body response tasks on humanoids. As next steps, we plan to deploy the prediction-free policy on hardware, move toward egocentric vision operation, and layer strategy learning via self-play so the robot can make its own hitting decisions.
@@ VII-B Training Details
S1 $r_{\text{global}}$ joint_position_limit -30
S1 $r_{\text{global}}$ joint_velocity_limit -0.1
S1 $r_{\text{global}}$ joint_torque_limit -0.1
@@
S2 $r_{\text{global}}$ joint_position_limit -30
S2 $r_{\text{global}}$ joint_velocity_limit -0.1
S2 $r_{\text{global}}$ joint_torque_limit -0.5
@@
S3 $r_{\text{loco}}$ momentum_positive 5
@@
S3 $r_{\text{loco}}$ face_the_net 8
@@
S3 $r_{\text{global}}$ joint_position_limit -30
S3 $r_{\text{global}}$ joint_velocity_limit -0.1
S3 $r_{\text{global}}$ joint_torque_limit -0.5

### v2

@@ IV-B1 Generate Shuttlecock Trajectory for Training
To generate badminton flight trajectories for training purposes, we employ a physics-based simulation approach. This method adheres to the badminton dynamics model in [30], with the shuttlecock’s flight state updated according to the following equation:
@@
where $m$ represents the shuttlecock mass, $v$ denotes its velocity, $g$ is gravitational acceleration, and $L$ is the aerodynamic characteristic length, defined as $L=2m/\rho SC_{D}$. Here, $\rho$ is air density, $S$ is the cross-sectional area of the shuttlecock, and $C_{D}$ is the drag coefficient. The computed aerodynamic length $L$ used in our work is 3.4.
The generated shuttlecock flight trajectories were required to meet specific conditions to ensure data validity and relevance. To simulate realistic badminton hitting scenarios, we defined a hitting zone constrained by $x\in[-0.8,0.8]$, $y\in[-1,0.2]$, and $z\in[1.5,1.6]$, with a minimum traversal time of 0.8 seconds. This filtering process effectively eliminated trajectories that did not conform to realistic hitting requirements, thereby ensuring the quality of the training data. Hitting zone in the $y$ direction is asymmetric due to the right-hand racket setting.
Finally, the selected trajectories were combined with corresponding interception point data to construct a comprehensive training dataset containing position, orientation (trajectory tangential line as the racket normal), and timing features. These data are stored as tensors and can be directly utilized for subsequent training.
@@ IV-B2 Shuttlecock Trajectory Prediction for Deployment
In the deployment phase, an EKF-based trajectory prediction algorithm was developed to achieve real-time shuttlecock flight estimation and hitting point prediction. This approach accounts for aerodynamic drag, gravitational acceleration, and sensor noise, providing accurate targets for the robot’s striking motion planning.
Our Extended Kalman Filter implementation adheres to the same badminton dynamics model as that used for shuttlecock trajectory generation. It operates through two distinct phases: measurement update and prediction. When the predicted height first falls into the predefined interception box, the spatial coordinates and time of traversal are designated as the predicted hitting target, expressed as: $\mathbf{p}_{\mathrm{ee}}^{*}=[x_{\mathrm{hit}},\,y_{\mathrm{hit}},\,z_{\mathrm{target}}]^{\mathsf{T}}$ and $t^{*}$. The corresponding velocity vector at $t_{hit}$ is then converted into a quaternion representation to guide the robot’s end-effector orientation during the hitting motion. The prediction is activated at 0.36 s, once sufficient trajectory history has been collected, and then continuously updates the hitting target in a rolling manner, which is fed to the controller.
@@ IV-C Prediction-free variant
We additionally train a prediction-free policy whose only change from the main pipeline lies in the observation space; the reward function, action space, training schedule (three-stage) remain identical. In this variant, the actor no longer receives the commanded hitting position, orientation, and time $(p_{ee}^{*},q_{ee}^{*},t^{*})$. Instead, it ingests a sliding window of world-frame shuttle positions comprising the current frame and the previous five frames sampled at 50 Hz (i.e., $\{p_{\mathrm{sh}}(t),p_{\mathrm{sh}}(t{-}0.02),\ldots,p_{\mathrm{sh}}(t{-}0.10)\}$). From this short history the actor must implicitly infer the intended interception pose and timing. The critic retains privileged access to the ”actual” interception point. This design is enabled by how training data are generated: for each commanded target, we not only sample the single interception point but also store the entire shuttle trajectory integrated by forward badminton dynamics. During training, the reward and the critic always know the ground-truth hit specification associated with the current shot, while the actor only observes the raw shuttle-history window.
A potential advantage of this variant is that deployment no longer relies on a hand-tuned predictor (e.g., an EKF plus a parametric aerodynamic model) to deliver $(p_{ee}^{*},q_{ee}^{*},t^{*})$. Instead, the controller conditions directly on measured shuttle positions, making the overall pipeline more end-to-end and reducing explicit dependence on predictor or aerodynamic parameters. During training, we randomize aerodynamic parameters per shot, effectively exposing the policy to an “aerodynamic patch” whose exact coefficients need not be known. Intuitively, this mirrors human play: players do not know a specific shuttle’s drag constants a priori but infer landing tendencies from a brief flight segment. In simulation, the prediction-free policy achieves hitting performance with only a modest drop relative to the known-target setting, providing initial evidence that our system can cope without an explicit shuttle predictor. We do not evaluate potential deployment benefits of this variant on hardware, real-robot validation and a more thorough analysis of its robustness are left to future work.
@@ V EXPERIMENTS
We evaluate our approach with five experiments that together probe accuracy, agility, robustness, and deployability: (1) a simulation Two-Robot Rally that measures sustained rally length; (2) a simulation Target-Known vs. Prediction-Free Comparison that contrasts two observation designs under matched conditions; (3) an EKF Prediction Accuracy study that quantifies trajectory and intercept-point errors; (4) a real-world Virtual-Target Swinging test that assesses swing speed and pose accuracy without a real shuttle; and (5) a Real-World Shuttle Hitting test that measures outgoing shuttle speed, reachable interception range, and return shot quality.
@@ V-B Simulation Results
Figure 3: Simulation results. Figure (a) illustrates the Two-Robot Rally scenario, where two identical humanoid robots sustain a rally of 21 consecutive returns. Figure (b) demonstrates the Prediction-Free policy: the robot infers the optimal impact position and orientation solely from the first five recorded shuttlecock positions after serving. Figure (c) presents the Target-Known policy, where a predetermined hitting position is provided. The red sphere indicates the designated hitting location, while the green sphere confirms successful impact execution by the robot.
@@ V-B1 Two-Robot Rally
Each policy executes twenty hits sampled randomly. We report two quantities: the position and orientation errors at impact and the executed swing speed. As shown in Fig. 4, explicit target information is not indispensable: the policy can infer both the hitting target and the timing on its own with modest performance drop.
Figure 4: Comparison between target-known and prediction-free policy. The top part of this figure shows the position error for both strategies. The middle section of the figure shows the orientation error comparison, the orientation corresponds to the normal direction of the racket face. The bottom part of the figure compares swing velocity.
@@ V-D Real Robot Deployment
In our real-world evaluation, we decompose the shuttlecock hitting task into two prerequisites, trajectory prediction and swing execution, followed by a final integrative test.
@@ V-D1 EKF Prediction Accuracy study
Prediction error refers to the deviation between the model’s predicted shuttlecock trajectory and the actual measured trajectory. Accurate trajectory prediction is critical for the robot to hit the shuttlecock at the right time and position. To quantify prediction errors, we collected 20 authentic badminton flight trajectories using a motion capture system. Partial segments of these trajectories were used as measurement input for the EKF. The EKF was then employed to generate predictions, which were subsequently compared against the ground-truth contact positions and timing. The comparative results are presented in Fig. 5. At 0.6 s before impact, the mean predicted position error falls below 100 mm, with the error bound sharply converging over time. By 0.3 s prior to contact, the mean position error has narrowed to 10 mm. Meanwhile, interception timing prediction achieves a mean error of approximately 20 ms at 0.6 s before hitting—already meeting task requirements—and similarly exhibits rapid convergence.
@@
To evaluate the impact of prediction model parameter uncertainty on prediction performance, a characteristic length perturbation factor was randomly generated for each trajectory to reflect changes in aerodynamic parameters; specific results are provided in the appendix.
@@ V-D2 Virtual-Target Swinging
The racket swing error refers to the spatial deviation between the racket’s contact point with the shuttlecock and the sweet spot during execution. To quantify the robot swing error, a large number of randomly sampled target positions were designated, and Mocap was employed to measure the center position of the racket face. The robot was subsequently instructed to execute the hitting motion, and the actual measured interception point was compared against the commanded target position, the corresponding results are presented in Fig. 6.
@@
We conducted a total of 71 virtual target swinging tests. The results show a mean Euclidean distance error of 24.29 mm with an average racket speed of 5.32 m/s at impact. These results verify that the robot can effectively control the contact point within the radius of the racket face during complex hitting tasks, confirming its practical feasibility for performing high-precision swinging actions.
@@ V-D3 Real-World Shuttle Hitting
When integrating two systems together in the real world, we set a ball machine to serve the shuttlecock and adjust its strength and pitch-yaw angles to test the capability of the policy. Because the shuttle’s relatively flat trajectories in the indoor mocap environment leave only a very short reaction window, we constrain the interception area to a reasonable range for safety. We collect 46 hitting results; except for 4 shots that hits the edge of racket frame, the robot successfully returns all shuttles (success rate 91.3%) with coordinated and agile motion. As shown in Fig. 7, the robot intercepts the shuttlecock at about 1.4-1.7 m above the ground with the region of 98 cm x 50 cm. This is a relatively large area for a robot height only 1.28 m. We also observe that the robot not only manages to intercept the shuttle but also returns it with high quality, achieving outgoing shuttle speeds of up to 19.1 m/s, with an average of 11.1 m/s across trials, and producing sharp slope trajectories whose peak height exceeds 3 m, as shown in Fig. LABEL:fig:figure1. Moreover, we measure the landing distance of the returned shots and obtain an average of about 4 m from the interception area, with only several outliers where the shuttle hits the ceiling or the racket frame.
Figure 7: Real-World Shuttle Hitting. This figure captures the robot’s actual hitting postures at the two opposing boundaries of the 98 cm × 50 cm interception area.
@@ VI-A System Design
We train a single whole-body policy that coordinates stepping and striking rather than layering an arm controller on top of a lower-body controller. In badminton, hit quality depends on how the legs reposition and support the arm’s exertion while maintaining balance. The subtle movements of the lower limbs can effectively adjust the posture of the racket.
Our multi-stage curriculum proceeds in three clear steps: learn to reach interception regions with stable footwork; add a sparse, high-value hit signal so a full swing emerges and follow a sigma curriculum to improve accuracy with light style terms; finally remove a majority of lower-limb regularizers to avoid gradients fighting the true objective of precise, energy-efficient hitting. During this stage, hitting metrics continue to improve while energy and torque costs drop, indicating progress toward task-optimal objective. Coupling position and orientation in the hit reward further prevents the policy from over-optimizing one at the expense of the other.
@@ VI-B Limitations and Future Work
Dependence on the mocap arena with limited ceiling height constrains shuttle trajectories to low arcs driven by a ball machine. During testing, shuttles occasionally hit the ceiling, and limited flight range makes long multi-ball rallies with human partners difficult. Higher ceilings and broader interception bands are expected to reveal more of the controller’s potential.
Moreover, the learned striking behavior remains relatively stereotyped. The current system predominantly discovers a single, forehand-like hitting style and does not yet exhibit the diverse repertoire of human strokes (e.g., backhand hits, lunges, jump and smashes), and the feasible interception area is restricted to a relatively narrow band.
Deploying the prediction-free variant on hardware remains future work. This variant removes explicit target prediction at deployment but relies on a strictly timed 50 Hz history of shuttle positions. Robust buffering and time-stamping will be needed so that the actor can infer velocities and timing from consistently spaced observations.
Moving toward pure vision would require reliable visual odometry and policies that keep the shuttle within the field of view during aggressive motion. A head-mounted camera that actively tracks the shuttle, together with learning signals that encourage aligning the view with the shuttle trajectory, appears to be a practical path to removing external tracking.
Above the low-level controller, a natural extension is to train a higher-level policy via multi-agent training to decide where to intercept, when to swing, and what racket-face orientation to command, and then invoke the whole-body controller as a motor primitive. Current failures arise when consecutive targets are too far apart for the available maneuver time; wider target sampling and reward shaping that explicitly encourages high-speed legged maneuvers are likely required to mitigate these cases.
Finally, the same framework may provide a path to adapt the pipeline to tennis or squash by replacing the shuttlecock dynamic “patch” and adjusting interception style, with modest reward retuning to reflect different sport-specific behaviors.
@@ VII Conclusions
In this work, we present the first real-world humanoid badminton system powered by a unified whole-body RL policy trained with a three-stage curriculum. The curriculum progresses from guiding lower-body footwork, to inducing a full swing and gradually tightening positional and orientation accuracy, and finally removing numerous gait regularizers to focus learning on precise, robust hitting. The policy produces human-like badminton behaviors that coordinate stepping and swinging under tight time windows, without any expert demonstrations.
In simulation, two identical humanoids sustain rallies with a best run of 21 consecutive returns, achieving a hitting accuracy within 0.10 m position error and 0.2 rad orientation error. A separate comparison shows that the prediction-free variant matches the behavior of its counterpart that uses known hitting targets, while remaining robust to moderate aerodynamic variation. On hardware, the controller achieves swing speeds of about 5.3 m/s and returns ball-machine serves with outgoing shuttle speeds up to 19.1 m/s under sub-second timing, transferring zero-shot through domain randomization. We believe this framework marks an important step toward agile, interactive whole-body response tasks on humanoids. As next steps, we plan to deploy the prediction-free policy on hardware, move toward egocentric vision operation, and layer strategy learning via self-play so the robot can make its own hitting decisions.
@@ VII-B Training Details
S1 $r_{\text{global}}$ dof_pos_limit -30
S1 $r_{\text{global}}$ dof_vel_limit -0.1
S1 $r_{\text{global}}$ dof_torque_limit -0.1
@@
S2 $r_{\text{global}}$ dof_pos_limit -30
S2 $r_{\text{global}}$ dof_vel_limit -0.1
S2 $r_{\text{global}}$ dof_torque_limit -0.1
@@
S2 $r_{\text{loco}}$ feet_no_fly -2
@@
S3 $r_{\text{loco}}$ face_the_net 5
@@
S3 $r_{\text{global}}$ dof_pos_limit -30
S3 $r_{\text{global}}$ dof_vel_limit -0.1
S3 $r_{\text{global}}$ dof_torque_limit -0.5